# Boosting CLIP Adaptation for Image Quality Assessment via Meta-Prompt Learning and Gradient Regularization

### v1

@@ 4.6 Ablation Study
Effect of Meta-Prompt Pre-training Module. This module consists of three key components: meta-learning, text meta-prompts, and visual meta-prompts. In Tab. 4, we perform ablation studies to evaluate the zero-shot capabilities of these components across various datasets. The baseline uses a CLIP model pre-trained on classification tasks. The second row shows meta-learning without prompts, fine-tuning CLIP’s visual and text encoders. The third and fourth rows focus on fine-tuning only the text or visual prompts, respectively, during meta-learning. The fifth row presents prompt tuning without meta-learning. Specifically, the second row reveals that meta-learning allows CLIP to learn shared quality knowledge across different distortions, improving IQA performance. However, adjusting CLIP’s weights reduces its original generalization, leading to lower performance on KonIQ. The third and fourth rows show that fine-tuning only the text or visual prompts significantly enhances zero-shot testing, especially on real-world datasets like KonIQ and LIVEC, by preserving CLIP’s generalization. The best results occur when both prompts are used together, confirming their mutual benefit. Finally, comparing the last row with the fifth shows that without meta-learning, prompt tuning leads to overfitting and reduced generalization.
Figure 4: (a) SRCC value during training. (b) Angles between $\bm{G}_{qua}$ and $\bm{G}_{sem}$ during training on KonIQ dataset. Under the constraints of QGR, our GRMP-IQA effectively captures the correlation between downstream quality knowledge and upstream general semantic knowledge.
@@
Effect of Meta-learning. To further investigate whether the effectiveness of our method derives from meta-learning, we conducted an ablation study. Specifically, as shown in rows 1 and 2 of Tab. 5, when using Empirical Risk Minimization without meta-learning pre-training on two synthetic dataset, the fine-tuning performance on LIVEC and KonIQ dataset was even worse than the baseline without pre-training. This indicates that additional pre-training data does not necessarily enhance performance and may even lead to overfitting. In contrast, the proposed meta-learning strategy allows the model to effectively leverage the generalizable quality priors obtained from the available external data, leading to significant performance improvements across different datasets.
Effect of Gradient Regularization. We performed ablation studies on the Quality-Aware Gradient Regularization (QGR) module. As shown in Tab. 6 and Fig. 4(a), without QGR, especially with very limited training data (e.g., 50 images), the model is prone to overfitting, leading to a sharp drop in generalization performance. QGR, by modulating training gradients, effectively reduces this overfitting and significantly improves adaptability to the test dataset.
@@
To illustrate QGR’s impact, we analyzed the angular difference between gradients $G_{qua}$ and $G_{sem}$ during training, as shown in Fig. 4(b). Without QGR, the angle between $G_{qua}$ and $G_{sem}$ tends toward 90 degrees, reflecting orthogonality typical of high-dimensional random vectors (Cai, Fan, and Jiang 2013). In contrast, QGR fosters a correlation between quality and semantic directions, improving the model’s ability to process quality-related information and reducing overfitting in limited datasets.
Effect of soft weight $\lambda$. As detailed in Tab. 7, a small $\lambda$ weakens QGR’s effectiveness, while a large $\lambda$ causes significant gradient changes and reduces performance. Based on this trade-off, we adopt $\lambda=5$ in our experiments.
@@
Figure 5: gMAD results between DEIQT (Qin et al. 2023) and ours. (a) Fixed DEIQT at low quality. (b) Fixed DEIQT at high quality. (c) Fixed ours at low quality. (d) Fixed ours at high quality.
5 Qualitative Analysis
To further assess our framework’s generalization, we trained models on the entire LIVE database and then tested them using the gMAD competition (Ma et al. 2016b) on the Waterloo Database (Ma et al. 2016a). gMAD efficiently selects image pairs with maximum quality difference predicted by an attacking IQA model to challenge another defending model which considers them to be of the same level of quality. The selected pairs are shown to the observer to determine whether the attacker or the defender is robust. As shown in Fig. 5, In the first two columns, our model attacks the competing method DEIQT, where each column represents images chosen from the poorer and better quality levels predicted by the defender. In the last two columns, we fix our model as the defender, giving image pairs selected from poorer and better quality levels, respectively. From Fig. 5, it is evident that when our model serves as the defender, the image pairs chosen by the attacker show little perceptual quality change, whereas, as the attacker, our model selects image pairs with more significant quality differences in succession. This indicates that the model has strong defensive and offensive capabilities. Additionally, it is important to highlight that the image pairs in the second column, which share similar semantic information, misled the DEIQT into classifying them as similar quality. Conversely, our model effectively identified the quality differences between them. These findings further underscore the strong generalization capability of our model in tackling complex distortions in real-world images.
@@ 6 Conclusion
In this research, we introduce the Gradient-Regulated Meta-Prompt Image Quality Assessment (GRMP-IQA) framework, which excels in generalization with limited data samples. This framework features a meta-learned soft prompt initialization module that quickly adapts to specific IQA tasks using pre-trained meta-knowledge. Additionally, an adaptive gradient regulation module refines the gradient trajectory during fine-tuning, focusing updates on image quality assessment while minimizing the impact of semantic content. Comprehensive experiments on various BIQA datasets validate the superior generalization of our framework, especially in data-scarce scenarios.

### v2

@@ 4.5 Ablation Study
Effect of Meta-Prompt Pre-training Module. This module consists of three key components: meta-learning, text meta-prompts, and visual meta-prompts. We conduct ablation studies in Tab. 5 to evaluate their zero-shot capabilities across various datasets. The baseline uses a CLIP model pre-trained on classification tasks. Row 2 evaluates meta-learning without prompts, where CLIP’s visual and text encoders are fine-tuned to learn distortion knowledge across different distortions, improving IQA performance. However, tuning CLIP’s weights impairs its original generalization, resulting in lower performance on KonIQ. Rows 3 and 4 assess fine-tuning only text or visual prompts during meta-learning. These strategies notably boost zero-shot performance—particularly on real-world datasets like KonIQ and LIVEC—by preserving CLIP’s generalization while adapting to IQA tasks. Row 5 explores prompt tuning without meta-learning, underscoring meta-learning’s role in curbing overfitting and maintaining generalization. The strongest results emerge when all components are combined, highlighting their complementary strengths.
Figure 4: (a) SRCC value during training with 50 samples. (b) Angles between $\bm{G}_{qua}$ and $\bm{G}_{sem}$ during training on KonIQ. Under the constraints of QGR, our method captures more correlation between quality knowledge and general semantic knowledge.
@@
Effect of Meta-learning. To further investigate whether the effectiveness of our method derives from meta-learning, we conducted an ablation study. Specifically, as shown in rows 1 and 2 of Tab. 4, when using Empirical Risk Minimization without meta-learning pre-training on two synthetic dataset, the fine-tuning performance on LIVEC and KonIQ dataset was even worse than the baseline without pre-training. This indicates that additional pre-training data does not necessarily enhance performance and may even lead to overfitting [19]. In contrast, our meta-learning strategy effectively extracts generalizable quality priors from external data, significantly enhancing performance across datasets.
@@
Effect of Gradient Regularization. We conducted ablation studies on the QGR module. As shown in Tab. 6 and Fig. 4(a), without QGR, models are prone to overfitting under limited training data, with rapid performance saturation and a slight decline, which ultimately hinders generalization. By modulating training gradients, QGR effectively mitigates overfitting and enhances adaptability to the test dataset. To illustrate QGR’s impact, we analyzed the angular difference between gradients $\bm{G}_{\text{qua}}$ and $\bm{G}_{\text{sem}}$ during training, as shown in Fig. 4(b). Without QGR, the angle between $\bm{G}_{\text{qua}}$ and $\bm{G}_{\text{sem}}$ tends toward 90 degrees, reflecting the orthogonality typical of high-dimensional random vectors [2]. In contrast, QGR introduces greater angular variation, suggesting that it encourages the model to explore more correlations between quality and semantic directions. This improves the model’s ability to capture quality-related information and mitigates overfitting on limited data.
Effect of soft weight $\lambda$. On the KonIQ dataset, we conduct ablation experiments with various soft weights $\lambda$ in Eq. 8 to examine their impact, as detailed in Tab. 7. Results show that a small $\lambda$ diminishes the effectiveness of our QGR, whereas a large $\lambda$ causes substantial gradient changes and decreases performance. Given our observation of the trade-off, we adopt $\lambda=5$ in our experiments.
@@
Figure 5: gMAD results between DEIQT [33] and our GRMP-IQA. (a) Fixed DEIQT at low quality. (b) Fixed DEIQT at high quality. (c) Fixed ours at low quality. (d) Fixed ours at high quality.
Method Param Efficiency Performance
Throughput ↑ Latency ↓ 80% Train 20% Train
Q-Align 8.2B 9.9 (img/sec) 101.0 (ms) 0.941 / 0.940 0.901 / 0.903
Ours 151M 43.0 (img/sec) 21.3 (ms) 0.945 / 0.934 0.931 / 0.915
Table 8: Comparison with Q-Align across varying training data ratio to evaluate efficiency and PLCC/SRCC on the KonIQ dataset.
Computational Analysis. As shown in Tab. 8, we evaluate the inference latency and throughput of our method on an RTX3090 GPU, surpassing Q-Align [45] in both efficiency and accuracy. Notably, our model has only 41M learnable and 151M total parameters, far fewer than Q-Align’s 8.2B.
4.6 Qualitative Analysis
The gMAD competition [28] is a standard method for assessing IQA model robustness. It selects image pairs where the attacker model predicts a large quality difference, while the defender perceives them as similar. Observers then evaluate these pairs to gauge model generalization. As shown in Fig. 5, when our model acts as the defender, the attacker’s selected pairs exhibit minimal perceptual quality changes. Conversely, as the attacker, our model consistently identifies pairs with significant quality differences. This highlights its strong defensive and offensive capabilities. Notably, in the second column, DEIQT misclassifies semantically similar images as having comparable quality, whereas our model accurately distinguishes their differences.
@@ 5 Conclusion
In this paper, we propose the GRMP-IQA framework, which generalizes well with limited data. It includes a meta-learning pre-training module that enables the CLIP model to rapidly adapt to IQA tasks and an adaptive gradient regulation module that refines gradient trajectories during fine-tuning, focusing updates on quality-aware knowledge while minimizing the negative impact of over-reliance on semantic noise. Comprehensive experiments on various BIQA datasets validate the superior generalization of our framework, especially in data-scarce scenarios.
Acknowledgements. This work was supported by the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. U23A20383, No. U21A20472, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305 and No. 62272401), and the Natural Science Foundation of Fujian Province of China (No. 2021J06003, No.2022J06001).